f(x)=x^2+1 find for f(x+2)
f(x+2)=(x+2)^2 +1
=x^2+4+1
=x^2 +5
But book answer is: f(x+2)=x^2+4x=5
I can not figure out where you get the 4x- what step am I missing?

Answer by nerdybill(7384)

You have to apply FOIL
.
f(x)=x^2+1 find for f(x+2)
.
f(x+2)=(x+2)^2+1
f(x+2)=(x+2)(x+2)+1
f(x+2)=(x^2+2x+2x+4)+1
f(x+2)=(x^2+4x+4)+1
f(x+2)= x^2+4x+5